# Non-Collaborative User Simulators for Tool Agents

### v2

@@ 4.2 Main Results
In the unavailable service mode, users request services beyond agents’ available APIs, requiring agents to decline or ignore these requests. Our results show consistent performance degradation across all models. When the agents encounter unavailable services from the user, they tend to repeatedly call the same helper APIs—functions that retrieve API documentation—leading to redundant calls in MultiWOZ (Table 9). This suggests that agents struggle to find solutions, repetitively re-fetching already-loaded API documentation. This helper API repetition appears to be a primary mechanism for performance degradation. GPT-4.1-nano’s failure rate until reaching max reasoning limit increases significantly in MultiWOZ when comparing the unavailable service mode to the collaborative mode, indicating that redundant helper API calls consume reasoning turns and prevent the agent from declining unavailable services within the given limit (Table 13). Interestingly, Qwen models exhibit consistently low duplication rates across both Qwen3-30b-a3b and Qwen3-235b-a22b, avoiding the helper API repetition problem. However, their performance outcomes differ dramatically: Qwen3-30b-a3b maintains minimal performance drop (97.7% relative SR), while Qwen3-235b-a22b shows substantial degradation (30.2% relative SR). The reason for this divergence lies in an alternative failure mechanism: Qwen3-235b-a22b shows a sharp increase in API result hallucinations—fabricating API results rather than obtaining them from actual calls—despite avoiding helper API repetition (Table 10).
@@
Figure 4: SFT Trained model comparison on MultiWOZ.
Figure 5: Human evaluation between PBUS and our user simulator
@@
In practical deployments, organizations often fine-tune smaller LLMs on their specific task data to reduce computational costs while maintaining performance. We investigate how small LLMs trained exclusively on collaborative user data perform when encountering non-collaborative behaviors. We fine-tuned three smaller models (Llama-3.2-3b-instruct, Qwen2.5-3b-instruct, Qwen2.5-7b-instruct) on successful dialogue completions between GPT-4.1-mini and a collaborative user simulator in MultiWOZ. As shown in Figure 5, fine-tuning enables smaller models to achieve over 90% success rates on collaborative users, which is a significant gain over the zero-shot setting (mostly below 4%). However, performance gains on non-collaborative users lag behind, particularly for the unavailable service and incomplete utterance modes. Fine-tuned models exhibit higher rates of duplicated helper API calls in the unavailable service mode and API input parameter hallucination in the incomplete utterance mode compared to zero-shot baselines (compare Table 15 with Table 1). These results demonstrate that fine-tuning on collaborative data alone produces agents with vulnerability to non-collaborative user behaviors, highlighting the need for diverse training strategies. We provide details about fine-tuning in §C.5.
@@
Real-world users often exhibit multiple non-collaborative behaviors within a single dialogue. Table 2 (Ours) shows GPT-4.1-mini’s performance when encountering two non-collaborative behaviors simultaneously. Notably, even GPT-4.1-mini, which demonstrated minimal performance degradation with individual non-collaborative behaviors (Table 1), experiences significant performance drops when facing users exhibiting two behaviors concurrently. These results show that performance degradation becomes more pronounced when models face multiple non-collaborative behaviors simultaneously.
@@ 4.3 User Simulator Evaluation
In this section, we evaluate our user simulator. An effective non-collaborative user simulator should model diverse user behaviors that deviate from collaborative patterns. At the same time, it must ensure that user utterances remain natural and provide all necessary information in the user goal for task completion. While §4.2 evaluated agent performance with single non-collaborative behaviors, we now examine more rigorous scenarios by combining multiple behaviors simultaneously. We compare our approach against the Prompt-Based User Simulator (PBUS) baseline, which follows a prompt-only approach used in $\tau$-bench (Yao et al., 2024). Unlike our user simulator, it has no additional LLM modules and solely adds non-collaborative behavior descriptions to the prompt. Through this comparison, we demonstrate that our user simulator architecture excels over prompt-only user simulators in conducting challenging simulations through the combination of multiple behaviors while ensuring goal alignment and dialogue quality.
@@
To verify the realism of our simulator’s non-collaborative behaviors, we conducted pairwise human evaluation with nine annotators evaluating 180 dialogues. For each dialogue pair, evaluators answered three questions: two assessing the realism of the individual behaviors and one evaluating that of the combined behaviors (see §D for detailed questionnaire and experimental setup). Figure 5 shows mixed results across different behaviors. Notably, our simulator significantly outperformed PBUS in incomplete utterance mode, as PBUS struggles to generate such behaviors through prompting alone. However, our simulator underperformed in tangential mode, primarily due to the tangential + incomplete utterance combination, where our simulator’s truncated tangential responses made the intended content less apparent to evaluators. When assessing combined behaviors, our simulator achieved approximately 60% win rate. These results confirm our simulator generates realistic non-collaborative behaviors comparable to or exceeding PBUS, demonstrating that the observed performance degradation stems from effective implementation of challenging behaviors rather than unrealistic dialogue generation.
@@ 5 Conclusion
Appendix A Collaborative User Simulator
A.1 Collaborative User Prompt
@@
A.2 Dialogue State Tracker
@@
The format of information pieces used in the dialogue state tracker differs between MultiWOZ and $\tau$-bench. In MultiWOZ, user goals exist in a structured JSON format, which is processed and used in the following format. In contrast, $\tau$-bench does not have user goals in JSON format, so we directly adopted the method from Laban et al. (2025) to shard user queries into multiple pieces of information. Table 3 shows examples of information pieces derived from user goals in MultiWOZ and $\tau$-bench.
@@ A.3 Ending Verifier
Appendix B Non-collaborative User Simulator
B.1 Unavailable Service
@@ B.2 Tangential
We define four tangential dialogue acts motivated by open-domain dialogue studies Yu & Yu (2021) as follows:
@@ B.3 Impatience
We define three impatient dialogue acts referring Reynolds & Harris (2009) and Harris & Reynolds (2004) as follows:
@@ B.4 Incomplete Utterance
Appendix C Experiments Details
C.1 Evaluation Environment Implementation Details
@@
We used the database from the MultiWOZ 2.4 dataset repository as is, and implemented the API ourselves in Python. We designed each domain to have one retrieve API and one book API, except for the taxi domain which only has a book API. We also included a “book_cancel” API. Additionally, we provided helper APIs from Trivedi et al. (2024), allowing the tool agent to utilize a total of 11 APIs during dialogue simulation. The tool agent uses the helper API as shown in Figure 6. First, it queries the available app (domain) descriptions using the “show_app_description” API. Then, it selects the appropriate app based on the context and queries the API description using “show_api_description”. Finally, the full documentation of the selected API is retrieved through “show_api_documentation”. In MultiWOZ, we commonly give all tool agents the following instruction prompt shown in Listing 17.
@@
We use the existing work’s environment with almost no modifications, and the only difference is the removal of the “transfer_to_human” API. Full implementation such as APIs, database and tool agent prompts of $\tau$-bench are in the Appendix of Yao et al. (2024)
C.2 Max Reasoning Limit
@@
C.3 Goal Alignment Metric
@@
C.4 Analysis Details
@@
C.5 Tool Agent Fine-tuning Details

### v3

@@ 4.2 Main Results
In the unavailable service mode, users request services beyond agent’s available APIs, requiring agents to decline or ignore these requests. Our results show consistent performance degradation across all models. When the agents encounter unavailable services from the user, they tend to repeatedly call the same helper APIs—functions that retrieve API documentation—leading to redundant calls in MultiWOZ (Table 10). This suggests that agents struggle to find solutions, repetitively re-fetching already-loaded API documentation. This helper API repetition appears to be a primary mechanism for performance degradation. GPT-4.1-nano’s failure rate until reaching max reasoning limit increases significantly in MultiWOZ when comparing the unavailable service mode to the collaborative mode, indicating that redundant helper API calls consume reasoning turns and prevent the agent from declining unavailable services within the given limit (Table 14). Interestingly, Qwen models exhibit consistently low duplication rates across both Qwen3-30b-a3b and Qwen3-235b-a22b, avoiding the helper API repetition problem. However, their performance outcomes differ dramatically: Qwen3-30b-a3b maintains minimal performance drop (97.7% relative SR), while Qwen3-235b-a22b shows substantial degradation (30.2% relative SR). The reason for this divergence lies in an alternative failure mechanism: Qwen3-235b-a22b shows a sharp increase in API result hallucinations—fabricating API results rather than obtaining them from actual calls—despite avoiding helper API repetition (Table 11).
@@
Figure 4: SFT training with collaborative and non-collaborative user simulation
@@
In practical deployments, organizations often fine-tune smaller LLMs on their specific task data to reduce computational costs while maintaining performance. We investigate how small LLMs trained exclusively on collaborative user data perform when encountering non-collaborative behaviors. We fine-tuned three smaller models (Llama-3.2-3b-instruct, Qwen2.5-3b-instruct, Qwen2.5-7b-instruct) on successful dialogue completions between GPT-4.1-mini and a collaborative user simulator in MultiWOZ. As shown in Figure 4, fine-tuning enables smaller models to achieve over 90% success rates on collaborative users, which is a significant gain over the zero-shot setting (mostly below 4%). However, performance gains on non-collaborative users lag behind, particularly for the unavailable service and incomplete utterance modes. Fine-tuned models exhibit higher rates of duplicated helper API calls in the unavailable service mode and API input parameter hallucination in the incomplete utterance mode compared to zero-shot baselines (compare Table 19 with Table 1). These results demonstrate that fine-tuning on collaborative data alone produces agents vulnerable to non-collaborative user behaviors. We provide details about fine-tuning in §D.6. Motivated by these observations, we fine-tuned Qwen2.5-3b-instruct using non-collaborative user data. This approach yielded performance improvements across all four non-collaborative user modes compared to models trained solely on collaborative user data. We provide detailed experimental settings, results, and discussion relevant to practical deployment in §D.7.
@@
Real-world users often exhibit multiple non-collaborative behaviors within a single dialogue. Table 2 (Ours) shows GPT-4.1-mini’s performance when encountering two non-collaborative behaviors simultaneously. Notably, even GPT-4.1-mini, which demonstrated minimal performance degradation with individual non-collaborative behaviors (Table 1), experiences significant performance drops when facing users exhibiting two behaviors concurrently. These results show that performance degradation becomes more pronounced when models face multiple non-collaborative behaviors simultaneously (An analysis of non-collaborative user behavior co-occurrence patterns is in §D.8).
User Simulator Extensibility
To demonstrate the extensibility of our user simulator, we applied our non-collaborative user simulator to ColBench’s (colbench) backend programming domain and MINT’s (mint) HotpotQA, HumanEval, TheoremQA, and GSM8K tasks. Since ColBench does not use tools and MINT involves solving goals through collaboration with the user simulator, we applied logics for certain behaviors (Details and results are in §F). These results demonstrate that our non-collaborative user simulation framework generalizes effectively across diverse task domains and interaction paradigms.
@@ 4.3 User Simulator Evaluation
In this section, we evaluate our user simulator. An effective non-collaborative user simulator should model diverse user behaviors that deviate from collaborative patterns. At the same time, it must ensure that user utterances remain natural and provide all necessary information in the user goal for task completion. While §4.2 evaluated agent performance with single non-collaborative behaviors, we now examine more rigorous scenarios by combining multiple behaviors simultaneously. We compare our approach against the Prompt-Based User Simulator (PBUS) baseline, which follows a prompt-only approach used in $\tau$-bench (tau_bench). Unlike our user simulator, it has no additional LLM modules and solely adds non-collaborative behavior descriptions to the prompt. Through this comparison, we demonstrate that our user simulator architecture excels over prompt-only user simulators in conducting challenging simulations through the combination of multiple behaviors while ensuring goal alignment and dialogue quality.
@@
To verify the realism of our simulator’s non-collaborative behaviors, we conducted pairwise human evaluation with nine annotators, comparing our simulator against PBUS across both single and combined behavior modes (See §E for detailed questionnaire and experimental setup). Figures 5(a) and 5(b) present the results for single behavior mode and combined behavior mode, respectively. Our simulator achieves approximately 60% overall win rate in both evaluations. In Figure 5(a), tangential behavior shows somewhat lower scores in certain combinations. This stems from the unnaturalness of some behavior combinations in goal-oriented dialogue contexts—for instance, tangential behavior (seeking sustained attention) conflicts with impatience (desiring quick service), and incomplete utterance (arising from typed communication) does not naturally co-occur with tangential behavior in task-oriented settings (See details in §D.8). Notably, tangential behavior achieves strong performance (66.7%) when evaluated in isolation, confirming that the lower scores in combined settings stem from unnatural pairings rather than limitations in simulating tangential behavior itself. These results confirm that our simulator generates realistic non-collaborative behaviors exceeding PBUS.
(a) Multiple Behavior Mode
(b) Single Behavior Mode
Figure 5: Human evaluation between PBUS and our user simulator
@@ 5 Conclusion
Appendix A Non-Collaborative User Behavior Taxonomy
A.1 Research Grounding
Marketing Research
Given that our work addresses goal-oriented dialogues where LLM agents assist users in achieving specific objectives, we drew from marketing literature that examines customer-service interactions, as these share fundamental similarities with user-agent dialogues. From marketing literature, we investigated the following user types:
  • Physical abuse toward service employees (pyhsical).
  • Aggressive verbal behavior in response to service dissatisfaction (non_coll_unavailable; service_staff_encounter).
  • Crime, fraud, and regulatory violations (service_staff_encounter; random_emotion).
  • Customers with illegitimate complaints (illegitimate).
  • Unavailable service requests (non_coll_unavailable).
  • Retaliation behaviors (retaliation).
  • Rapport-seeking customers (rapport_0; rapport_1; employee_rapport).
  • Customers demanding constant attention (stress_burnout; attention_demand).
User-LLM Conversation Studies
Since tool agent interactions represent human-LLM conversations, we investigated user characteristics identified in this emerging domain. From user-LLM conversation studies, we investigated the following user types:
  • Truncated utterances (lmsys; wildchat).
  • Anger expression toward LLM agents (lmsys; wildchat; crolic2022blame; chatbot_negative).
  • Toxic content generation (lmsys; wildchat).
  • Underspecification (getlost_multiturn).
A.2 Selection and Clustering
Selection
Since our focus is on simulating non-collaborative users in conversational contexts, we excluded behaviors from marketing literature that require physical interaction or temporal progression beyond dialogue scope. We retained behaviors relevant to conversational dynamics. From marketing literature, we focused on:
  • Aggressive verbal behavior in response to service dissatisfaction.
  • Customers with illegitimate complaints.
  • Unavailable service requests.
  • Rapport-seeking customers.
  • Customers demanding constant attention.
From user-LLM conversation studies, we excluded toxic content (due to ethical considerations) and focused on:
  • Truncated utterances.
  • Anger expression toward LLM agents.
  • Underspecification.
Clustering
We systematically clustered these behaviors based on their underlying characteristics. Each of our four categories represents multiple user types from marketing and user-LLM conversation research, ensuring comprehensive coverage of real-world scenarios. This approach grounds our taxonomy in theory while maintaining a manageable scope for simulation purposes.
  • Unavailable Service: Combines “customers with illegitimate complaints” and “unavailable service requests”
  • Tangential: Merges “rapport-seeking customers” and “customers demanding constant attention”
  • Impatience: Integrates “aggressive verbal behavior in service dissatisfaction” and “anger expression toward LLM agents”
  • Incomplete Utterance: Unifies “truncated utterances” and “underspecification”
Appendix B Collaborative User Simulator
B.1 Collaborative User Prompt
@@
B.2 Dialogue State Tracker
@@
The format of information pieces used in the dialogue state tracker differs between MultiWOZ and $\tau$-bench. In MultiWOZ, user goals exist in a structured JSON format, which is processed and used in the following format. In contrast, $\tau$-bench does not have user goals in JSON format, so we directly adopted the method from getlost_multiturn to shard user queries into multiple pieces of information. Table 3 shows examples of information pieces derived from user goals in MultiWOZ and $\tau$-bench.
@@ B.3 Ending Verifier
Appendix C Non-collaborative User Simulator
C.1 Unavailable Service
@@ C.2 Tangential
We define four tangential dialogue acts motivated by open-domain dialogue studies midas_tangential as follows:
@@ C.3 Impatience
We define three impatient dialogue acts referring bellgerent_abuse and threat as follows:
@@ C.4 Incomplete Utterance
Appendix D Experiments Details
D.1 Evaluation Environment Implementation Details
@@
We used the database from the MultiWOZ 2.4 dataset repository as is, and implemented the API ourselves in Python. We designed each domain to have one retrieve API and one book API, except for the taxi domain which only has a book API. We also included a “book_cancel” API. Additionally, we provided helper APIs from appworld, allowing the tool agent to utilize a total of 11 APIs during dialogue simulation. The tool agent uses the helper API as shown in Figure 6. First, it queries the available app (domain) descriptions using the “show_app_description” API. Then, it selects the appropriate app based on the context and queries the API description using “show_api_description”. Finally, the full documentation of the selected API is retrieved through “show_api_documentation”. In MultiWOZ, we commonly give all tool agents the following instruction prompt shown in Listing 17.
@@
We use the existing work’s environment with almost no modifications, and the only difference is the removal of the “transfer_to_human” API. Full implementation such as APIs, database and tool agent prompts of $\tau$-bench are in the Appendix of tau_bench
D.2 Max Reasoning Limit
@@
For additional verification, we present experimental results on MultiWOZ by extending the limit to 50 steps in Table 9. The experiments were conducted on GPT-4.1-nano and Qwen3-30b-a3b, which showed the lowest performance in Table 1.
Table 9: Performance on 50 limit reasoning in MultiWOZ
Qwen3-30b-a3b 47.8 50.0 34.6 41.0 34.3
GPT-4.1-nano 22.5 16.0 12.6 19.4 16.6
Overall, while some cases show improved performance, others exhibit performance decreases. This suggests that the agent’s low performance is not simply due to insufficient reasoning steps. Additionally, the performance gap between collaborative and non-collaborative modes persists at comparable levels regardless of the step limit. For instance, Qwen3-30b-a3b shows no performance drop in unavailable service scenarios in both settings, and GPT-4.1-nano at the 30-step limit does not experience performance degradation in impatience scenarios, while at the 50-step limit it shows a small decrease. Although Qwen3-30b-a3b shows improved performance in incomplete utterance and tangential modes, a significant performance gap between collaborative and non-collaborative modes remains. Based on these observations, we conclude that the performance gap between collaborative and non-collaborative settings in our experiments was not underestimated due to reasoning step limitations.
D.3 Goal Alignment Metric
@@
Causes of Goal Misalignment
We run simulations in collaborative mode and in each of the four non-collaborative modes individually on $\tau$-Bench and MultiWOZ. Among the total of 1,230 dialogues, only 37 cases (3.0%) exhibited goal misalignment. The breakdown of causes are as follows:
  1. 1.
    18 out of 37 simulations (49%) were marked as goal-misaligned despite actually being aligned, because the alignment checker failed to capture certain expressions of goal statements. For instance, common cases included the user simulator mentioning “Wi-Fi” or “Visa”, and the alignment checker failing to associate them with “internet” or “credit card” in the goal statements. However, such errors were rare and random (e.g., only 2% of “Wi-Fi” mentions and 3% of “Visa” mentions were missed; in the other 98% and 97% of cases, the alignment checker succeeded). After regeneration, the final goal-aligned simulations faithfully conveyed all intended information.
  2. 2.
    10 out of 37 simulations (27%) were marked as goal-misaligned because the user simulator did not mention certain goal statements when doing so would lead to unnatural dialogue. A notable example occurs when $\tau$-Bench goals contain conditional statements such as “If that’s not available, book the earliest flight to Newark the next day”. If the condition is not met (e.g., the agent successfully books the initially requested ticket), the user simulator would not state the conditional goal statement, as that would be unnatural. We think that filtering such cases is reasonable because conditional goal statements are meant to capture situations where backup plans or alternative suggestions become relevant or expected. If anything, the regenerated, goal-aligned scenarios reflect this intention more faithfully.
  3. 3.
    The remaining 9 out of 37 simulations (24%) involved situational statements about the user that were not mentioned by the user simulator (e.g., “I’m getting into gaming”). This occurred mainly because the simulated conversations did not flow in a way that made these statements natural to express. Again, such statements in user goals signal scenarios where they would become contextually relevant. The final goal-aligned scenarios therefore better reflect this intention. Moreover, these scenarios are not necessarily easier than the goal-misaligned ones.
Human Correlation
To measure the reliability of the judge model for goal alignment, we sampled 40 dialogues from Tau-bench and MultiWOZ, including both collaborative and non-collaborative user simulations (20 goal-aligned dialogues and 20 goal-misaligned dialogues), and instructed a human annotator (graduate student) to evaluate whether all information specified in the user goal was provided, given the user goal and dialogue. We then measured the Matthews correlation coefficient (MCC) between the human judgments and the judge model’s judgments (the MCC is the Pearson correlation coefficient for binary annotations). The correlation was 0.77, which is generally considered strong (mcc_strong) . This result suggests that the judge model has strong agreement with human judgments.
D.4 Analysis Details
@@
D.5 Excluded Test cases
In our experiments, we excluded three types of tasks.
MultiWOZ: Retrieval Tasks
MultiWOZ’s API has two types in each domain (we will use “restaurant” as an example): (1) restaurant_retrieval: retrieves the list of restaurants that match the conditions. This API provides the restaurant’s ID and other features. (2) restaurant_booking: uses the restaurant ID to book the restaurant. The user goal of retrieval tasks is to obtain a list of restaurants matching certain conditions. To solve this, the agent only needs to call the retrieval API and provide the list through an utterance. On top of this, booking tasks go one step further: the agent informs the user about the restaurant list, the user selects one from the options, and the agent books that restaurant using the booking API, resulting in task success. Since successful booking requires successful retrieval, our evaluation is already covering the agent’s capability on retrieval tasks. The main reason we evaluated task success or failure based on booking completion, excluding retrieval-only scenarios, is the difficulty and inaccuracy involved in determining the success or failure of retrieval tasks on their own. For booking tasks, we can straightforwardly determine success if the DB state is successfully changed as expected. However, retrieval tasks do not involve DB state changes. In addition, it is difficult to identify success or failure based on whether all necessary constraints are set in the parameters of retrieval API calls, because sometimes the agent performs constraint-based filtering on its own based on the retrieved entries. Therefore, the assessment would require LLM involvement as a judge, which still tends to produce inaccurate judgments. Since successful booking requires successful retrieval tasks, our evaluation can be seen as a stricter evaluation setting.
MultiWOZ: Single Domain Tasks
Although multi-domain (cross-domain) scenarios were sufficient for evaluation, because multi-domain scenarios essentially consist of single-domain dialogues plus additional challenges due to the multi-domain nature, we have conducted additional experiments on single-domain tasks and present the merged score (143 single-domain + 89 multi-domain) in Table 15.
Table 15: The merged result of the MultiWOZ single-domain and cross-domain
GPT-4.1-mini 94.0 91.4 90.1 89.2 88.0
GPT-4.1-nano 39.1 32.0 23.3 37.1 26.0
Qwen3-235b-a22b 81.8 73.9 65.4 72.5 67.2
Qwen3-30b-a3b 60.7 59.7 40.7 54.3 46.0
Llama-3.1-70b-instruct 76.0 70.5 64.9 65.3 65.6
The merged results show overall performance improvements compared to Table 1. However, the performance decrease in non-collaborative settings compared to collaborative settings remains consistent. Additionally, models exhibit performance degradation patterns in non-collaborative versus collaborative settings that are similar to those shown in Table 1. For example, Qwen3-30b-a3b, there was almost no performance change in unavailable service mode, a small performance decrease in impatience mode, and substantial performance decreases in tangential and incomplete utterance mode—a trend identical to Table 1. Furthermore, GPT-4.1-nano, which showed no performance decrease in impatience mode in Table 1, also exhibits the smallest performance decrease in impatience mode in the merged results.
$\tau$-Bench: Transfer to Human Scenarios
In our previous experiments, we found that agents resort to human transfer too frequently, especially when encountering unavailable service and impatience modes. Therefore, we excluded the “human transfer” action and the 8 test scenarios in $\tau$-bench where it is the correct answer. We present the results of our previous experiments with 60 samples from $\tau$-bench, simulated 4 times each for a total of 240 evaluations:
Table 16: Number of Human Transfer
Collaborative Unavailable Service Impatience
GPT-4.1-mini 17 58 87
GLM-4.5 16 39 54
DeepSeek-Chat-v3.1 27 78 75
The disproportionately high rate of human transfer occurrences would risk producing experimental results limited to domains or environments where human transfer functionality exists, which would not reflect the general robustness of agents against non-collaborative users. Therefore, we decided to exclude human transfer scenarios to ensure our evaluation framework captures more generalizable agent capabilities. Additionally, we present the performance results for the 8 excluded cases, simulated 4 times each for a total of 32 simulations.
Table 17: Performance of 8 scenarios where human transfer is the ground truth on $\tau$-Bench
GPT-4.1-mini 93.8 71.9 81.3 93.8 90.6
GPT-4.1-nano 56.3 65.6 62.5 56.3 62.5
Qwen3-235b-a22b 68.8 75.0 62.5 78.1 71.9
Qwen3-30b-a3b 84.4 68.8 87.5 78.1 81.3
Llama-3.1-70b-instruct 71.9 71.9 75.0 75.0 62.5
D.6 Tool Agent Fine-tuning Details